OpenStudy (anonymous):

i need some help solving for x 2^(x+1) + 8^(x-2) = 24 answer is 3 using trial and error, however, i need the solution. thank you in advance!

OpenStudy (anonymous):

as 8 = 2^3 we can rewrite the expression as: \[2^{x+1}+2^{3(x-2)}=24\] Then taking \[\log _{2}\] of both sides we get: \[\log _{2}(x+1) + \log _{2}(3x-6) = \log_{2}24\]

OpenStudy (anonymous):

another thing, we are not allowed to use log...

OpenStudy (anonymous):

ahhhh. that is unfortunate

OpenStudy (anonymous):

24 = 16 + 8 24 = 2^4 + 2^3 2^(x+1)+2^(3x−6)=2^4 + 2^3

OpenStudy (anonymous):

x+1 = 4 3x-6 = 3

OpenStudy (anonymous):

Those both solve to x=3 and x = 3

OpenStudy (anonymous):

yup - good work

OpenStudy (anonymous):

got the answer already

OpenStudy (anonymous):

it's the same with TBates! thank you!